OpenStudy (anonymous):

The perimeter of a rectangle is 760 feet. Let x represent the width of the rectangle. Write a quadratic function for the area A of the rectangle in terms of its width.

OpenStudy (raden):

do u know what's the formula to find the perimeter of a rectangle

OpenStudy (anonymous):

by adding all sides?

OpenStudy (raden):

represented in a formula ....

OpenStudy (anonymous):

P=2L+2W

OpenStudy (raden):

yes, right u are now plug P=760, what u get

OpenStudy (anonymous):

760=2L+2W

OpenStudy (raden):

then simplify it, both sides divided by 2

OpenStudy (anonymous):

380=L+W

OpenStudy (raden):

yes... now from 380 = l+w, it similar by l=380-w right ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, now plug w=x gives l = 380-x, right ?

OpenStudy (anonymous):

yeah, I understand

OpenStudy (raden):

then, u know that the formula to find area of a rectangle is A = l * w subtitute l and w in x form above, what u get

OpenStudy (anonymous):

380x-x^2

OpenStudy (raden):

A = (380-x)x = 380x-x^2 u are right

OpenStudy (anonymous):

so as a quadratic would I write it as -x^2-380x

OpenStudy (anonymous):

-x^2+380x

OpenStudy (raden):

thought one :)